ADNAN MUHAMMED

Grade 12,

A body of mass m ,having momentum p is moving on a rough horizontal surface . If it stopped in a distance x ,the cofficient of friction between the body and the surface is ?

A body of mass m ,having momentum p is moving on a rough horizontal surface . If it stopped in a distance x ,the cofficient of friction between the body and the surface is ?
 

Grade:11

1 Answers

Amisha Porwal
6 years ago
Its velocity will be u=p/m
Now acc to  v2=u2-2as
                   (  P/m)2=2ax
                    a=p/m)2/2x
f=m*a=coeff *normal
coeff=(p/m)2/2gx